dy/dx=2x+2xy²/y+2x²y


1
Expert's answer
2020-10-18T17:48:02-0400

"\frac{dy}{dx}=\frac{2x+2xy^2}{y+2x^2y}"


Let us solve this differential equation:


"\frac{dy}{dx}=\frac{2x(1+y^2)}{y(1+2x^2)}"


"\frac{ydy}{1+y^2}=\frac{2xdx}{1+2x^2}" ("1+y^2>0")


"\int\frac{ydy}{1+y^2}=\int\frac{2xdx}{1+2x^2}"


"\frac{1}{2}\int\frac{d(1+y^2)}{1+y^2}=\frac{1}{2}\int\frac{d(1+2x^2)}{1+2x^2}"


"\int\frac{d(1+y^2)}{1+y^2}=\int\frac{d(1+2x^2)}{1+2x^2}"


"\ln(1+y^2)=\ln(1+2x^2)+ln(C)"


"\ln(1+y^2)=\ln((1+2x^2)C)"


Therefore, the general solution is the following:


"1+y^2=C(1+2x^2)"